# SAM 2++: Tracking Anything at Any Granularity

### v2

@@ 6.1 Comparison to state-of-the-art on three tasks
Video Object Segmentation. The comparisons between our model and previous semi-supervised VOS methods are demonstrated in Table 3, including YoutubeVOS-19 [60], MOSE [19], LVOS-v2 [29], BURST [1], VOST [51], VISOR [18], and our TAG dataset. We use the standard metric $\mathcal{J\&F}$ [47] that averages Jaccard index and contour accuracy in most benchmarks, but adopt Higher Order Tracking Accuracy (HOTA) [43] in the BURST benchmark. Results show that our model outperforms individual video object segmentation models.
Single Object Tracking. We compare the performance of our proposed model on three benchmarks in Table 4, including TrackingNet [44], GOT-10k [31], TNL2K [54], VastTrack [46] and our TAG dataset, and all compared models are trained on four datasets. We choose the Average Overlap (AO) for the GOT-10k benchmark, and Area Under the Curve (AUC) for the other benchmarks. Experiments demonstrate that our model consistently outperforms previous state-of-the-art SOT approaches across all benchmarks.
Table 5: State-of-the-art comparison on Point Tracking Task.
Methods BADJA PerceptionTestval $_{PT}^{test} TAG$$_{PT}^{val}$ Tapviddavis Tapvidrgb
pips [27] 64.2 41.5 19.0 19.8 40.9 28.5
pips [27] 56.9 27.2 12.3 12.2 26.3 20.8
pips++ [71] 56.6 62.6 20.9 23.1 59.8 58.5
CoTracker [34] 65.3 59.8 23.3 22.3 60.9 63.1
CoTracker [34] 55.1 48.9 18.8 18.1 50.8 46.1
CoTracker3 [33] 72.7 71.3 29.6 29.1 65.6 70.6
CoTracker3 [33] 66.3 66.3 25.8 24.9 59.2 63.6
TAPTR [38] 69.1 59.4 23.7 23.8 61.2 58.0
TAPTR [38] 63.0 48.9 20.4 19.0 52.3 39.4
TAPIR [22] 64.2 59.6 21.3 24.6 56.8 50.4
LocoTrack [15] 68.7 67.1 25.2 30.2 62.7 69.2
Track-On [2] 69.7 69.7 24.8 25.8 64.5 64.5
Ours 72.9 66.2 35.3 37.7 56.1 59.0
@@
Online Point Tracking. We compare our method to prior works in Table 5 on four benchmarks, including BADJA [5] and Perception Test [49] for key point tracking, TAP-Vid [21] for arbitrary point tracking, and our TAG dataset in the ‘query first’ evaluation, which means points appearing in the first frame are used as queries. We report the Percentage of Correct Keypoint-Transfer (PCK-T) for the BADJA benchmark, and Average Jaccard (AJ) for the remaining benchmarks. However, most of the current methods are offline trackers, which process long-temporal window frames or even the entire video to be able to see the future frame, and do not match the online setup required by real applications. For a fair comparison, we modified their input so that there is no future information inside the window, denoted as model, to enable inference online. Experimental results show a substantial decrease in model performance when the input data is switched from offline to online. The comparative analysis reveals the effectiveness of our approach on keypoint tracking benchmarks, which surpasses competing models. Furthermore, although our model is trained on keypoint datasets, it demonstrates generalization capability on arbitrary point tracking datasets.
@@ 6.2 Exploration Studies
Study on data engine. To validate the effectiveness of our data engine, we evaluated the performance when trained on different phases of our TAG dataset, as shown in Table 6. The results demonstrate that our proposed dataset enhances the performance on other datasets, indicating high diversity and generalizability. After training with data from more phases, the performance is further improved, demonstrating the effectiveness of the supplementary data provided by our data engine.
@@ 6.3 Visualization
To further illustrate the varying requirements for memory representation of targets at different granularities, we visualize the memory-related outputs for the three tasks under both task-adaptive memory mechanism (left) and full parameter sharing (right), as shown in Fig. 6. Firstly, we observe that even under different training settings, the memory features for the same task remain highly similar, indicating that different granularities have distinct memory requirements. Secondly, under the memory-related decoupled training setting with our task-adaptive memory mechanism, both memory attention and mask output align more closely with the task outputs compared to full parameter sharing, highlighting the necessity of the decoupled design. Finally, we find that under the full parameter sharing setting, the mask output for point tracking does not exhibit a Gaussian pattern, leading to incorrect predictions. This demonstrates that the decoupled design effectively preserves the specific needs of different tasks.
@@ 7 Conclusion
More importantly, further model results and dataset annotations are displayed on the https://tracking-any-granularity.github.io/, where our datasets and models will also be open-sourced.
@@ 8.2 Training and Inference Details
Training is performed jointly on data of the three tasks. In addition to our Tracking-Any-Granularity dataset, we used DAVIS-17 [48], YoutubeVOS-19 [60] and MOSE [19] for the mask task, LaSOT [23], GOT-10k [31], TrackingNet [44] and COCO [41] for the box task, and TAP-Vid Kinetics [21], PointOdyssey [71], and PerceptionTest [49] for the point task. To enable the model to be simultaneously capable of all three tasks and to optimize training efficiency, we adopt the strategy of alternating between the three tasks. Specifically, we implement parallelisation by sampling a whole batch at each step of training, which is entirely derived from the data of a particular task. The sampling probability is set to 1:4:5 to balance the performance of the three tasks.
@@
In addition, if interactive input is used as initial prompt, we select up to 2 frames as corrective frames to add corrective clicks on them: after predicting the selected frame, we sample a positive point from the false positive region between the prediction and ground truth or a negative point from the false negative region as a corrective point, and use it as additional prompt to get a new prediction along with all previous cumulative prompt from that frame. This operation is repeated until 7 corrective points have been added. In addition, if the box tracking task uses the box format to compute the regional differences between the prediction and the GT, there is an overwhelming problem that the sampled corrective clicks may fall at the boundaries of the box instead of inside the target, which is contrary to the actual interaction. Therefore, we choose to compute the difference in mask format, and use SAM 2 and sam-hq [35] with box annotations to obtain the pseudo-GT mask on SOT datasets because of its good segmentation ability.
Losses and optimization. Following the mask tracking task in SAM 2, we adopt the linear combination of focal loss $\mathcal{L}_{mask}^{focal}$ and dice loss $\mathcal{L}_{mask}^{dice}$ for the mask prediction, L1 loss for the IoU prediction $\mathcal{L}_{L1}^{IoU}$, and cross-entropy loss for object occlusion prediction $\mathcal{L}_{CE}^{obj}$. During the box tracking task, we adopt the corner head to predict the bounding boxes and add additional ciou loss [72] and L1 loss to supervise the box prediction. As for the point tracking task, we select the highest probability position from the mask prediction as point prediction, and use soft argmax [25] during training for making the process derivable instead of the undifferentiable argmax function. Beyond the loss on mask in the form of Gaussian map, we add an L1 loss between the prediction and ground-truth point to directly optimize the distance and accuracy of the points. For multi-prediction scenario, we only supervise the task predictions (masks, boxes, and points) with the lowest loss, which is a combination of $\mathcal{L}^{mask}$, $\mathcal{L}^{box}$ and $\mathcal{L}^{point}$, but supervise the IoU predictions of all task predictions to learn to synchronise the quality of predictions. Furthermore, if the target is missing in some frames due to disappearance or cropping, we do not supervise the task predictions or iou predictions on them in all three tasks, but always supervise the occlusion prediction from an MLP head, no matter if the ground-truth exists or not. In summary, the supervision losses for the three tasks can be written as:
@@ 9.2 Annotation Pipeline
3) Fine Annotation. To reduce annotator workload and improve efficiency, we use SAM [37] to generate rough masks based on the coarse annotations (points and boxes). Then, annotators refine these masks with the following requirements:
@@ 10.1 Performance Comparison
Evaluation metrics. In video object segmentation task, we use standard metrics [47] in most benchmarks: Jaccard index $\mathcal{J}$, contour accuracy $\mathcal{F}$, and their average $\mathcal{J\&F}$. In the YouTubeVOS benchmark, $\mathcal{J}$ and $\mathcal{F}$ are computed for ”seen” and ”unseen” categories separately. $\mathcal{G}$ is the averaged $\mathcal{J\&F}$ for both seen and unseen classes. In LVOS benchmark, the first densely annotated long-term VOS dataset with high-quality annotations, it introduces the standard deviation $\mathcal{V}$ of the average score of $\mathcal{J}$ and $\mathcal{F}$ to assess the temporal stability of VOS models. In VOST benchmark, which focuses on segmenting objects as they undergo complex transformations, it additionally reports $\mathcal{J}_{tr}$ for the last 25% of the frames in a sequence to show the robustness after the transformation has been mostly completed. The BURST benchmark is evaluated with Higher Order Tracking Accuracy (HOTA) [43] as a good balance between measuring frame-level detection and temporal association accuracy. In single object tracking task, we evaluate performance with Area Under the Curve (AUC), normalized precision (${P}_{Norm}$) and precision (P{P}) to measure the average accuracy of center, size and scale between the prediction and labeled groundtruth bounding boxes of all the frames for most benchmarks. For the GOT-10k benchmark, we choose the average overlap (AO) and success rate (SR) as indicators. The former AO denotes the average of overlaps between all groundtruth and estimated bounding boxes, while the SR measures the percentage of successfully tracked frames where the overlaps exceed a threshold (e.g., 0.5). In point tracking task, we report Occlusion Accuracy (OA) and Average Jaccard (AJ) for TAP-Vid [21], Perception Test [49], and our dataset. As for the BADJA benchmark, we adopt the Percentage of Correct Keypoint-Transfer (PCK-T). We measure these benchmarks in the ‘query first’ evaluation, which means points appearing in the first frame are used as queries.
@@
Unicorn [63] 69.2 65.2 73.2 83.0 86.4 82.2
UNINEXT-H [64] 81.8 77.7 85.8 85.4 89.0 86.4
MITS [61] 84.9 82.0 87.7 83.4 88.9 84.6
OmniTracker-L [53] 71.0 66.8 75.2 83.4 86.7 82.3
Ours 89.1 86.3 91.9 86.0 90.1 87.3
@@
Comparison with TAG model. We found that there exists a work, TAG [26], that shares the same objective as ours, which is to achieve tracking of mask, box, and point with a unified model. To highlight the contribution of our work, we provide a detailed comparison between the two approaches. Firstly, TAG is an offline tracking model that processes multiple frames as a clip simultaneously, which not only differs from the current mainstream online frame-by-frame tracking pipeline but also leads to information leakage from future frames and is only applicable to pre-recorded videos. In contrast, our model ensures that the current frame only receives information from the past frames, making it suitable for video streams. Secondly, in the way of the prompt construction, TAG simply converts point coordinates into a {0, 1} mask in the point task, providing limited target information. Our method combines point coordinates with a (0, 1) Gaussian mask: the former provides precise locations, while the latter highlights the target point in mask form, maintaining consistency with output from MaskDecoder and input for MemoryEncoder, thereby enhancing expressiveness. For the box task, TAG converts the box into a square mask, which causes confusion between the target region and the background, affecting the accuracy of the target information. Third, TAG is trained only on public datasets, which limits the scale of the training dataset. In contrast, we construct a Data Engine that enables both model training and dataset annotation expansion, ultimately resulting in a large-scale dataset with three types of granularity annotations and a well-trained model. Most important of all, as an offline approach, the TAG model primarily focuses on how to jointly encode targets of varying granularity. When processing the next clip, the prompt remains in an original, unmodeled form, lacking rich target representation (e.g., mask, point, or box). Meanwhile, due to the lack of judgment of predictions, the next clip must adopt the prediction of the last frame in the previous clip as a prompt, even if it may be unreliable, which leads to error accumulation and makes it difficult to handle common challenges such as temporary target disappearance. In contrast, our method follows the online setting. The core challenges lie not only in multi-granularity prompts encoding, but also in how to transform predictions into memory representations to guide subsequent frames. Compared to original prompts, memory offers richer target features that improve the stability and accuracy. Leveraging both selective capability and memory diversity, the mechanism compensates for potential errors in individual predictions, effectively improving stability. To make this key component compatible with multi-granularity tracking, we introduced a Task-Adaptive Memory mechanism, one of the major contributions, to unify the predictions of varying granularity into memory representations. In summary, the distinction is not merely at the task level, but it directly impacts motivation and core innovations. Additionally, we present a performance comparison on the three tasks. As shown in the Table 14, our model significantly outperforms the TAG model on all three tasks, demonstrating the superior performance of our model.

### v3

@@ 6.1 Comparison to state-of-the-art on three tasks
Video Object Segmentation. The comparisons between our model and previous semi-supervised VOS methods are demonstrated in Table 3, including YoutubeVOS-19 [youtube-vos], MOSE [MOSE], LVOS-v2 [LVOS_V2], BURST [burst], VOST [vost], VISOR [visor], and our TAG dataset. We use the standard metric $\mathcal{J\&F}$ [vos-benchmark] that averages Jaccard index and contour accuracy in most benchmarks, but adopt Higher Order Tracking Accuracy (HOTA) [hota] in the BURST benchmark. Results show that our model outperforms individual video object segmentation models.
Single Object Tracking. We compare the performance of our proposed model on three benchmarks in Table 4, including TrackingNet [TrackingNet], GOT-10k [got10k], TNL2K [wang2021tnl2k], VastTrack [vasttrack] and our TAG dataset, and all compared models are trained on four datasets. We choose the Average Overlap (AO) for the GOT-10k benchmark, and Area Under the Curve (AUC) for the other benchmarks. Experiments demonstrate that our model consistently outperforms existing methods across nearly all benchmarks.
Table 5: State-of-the-art comparison on Point Tracking Task. model and model means offline and modified online trackers.
Method Key Point Arbitrary Point
BADJA PerceptionTestval $_{PT}^{test} TAG$$_{PT}^{val}$ Tapviddavis Tapvidrgb
pips [harley2022pips] 64.2 41.5 19.0 19.8 40.9 28.5
pips [harley2022pips] 56.9 27.2 12.3 12.2 26.3 20.8
pips++ [zheng2023pointodyssey] 56.6 62.6 20.9 23.1 59.8 58.5
CoTracker [cotrackerv1] 65.3 59.8 23.3 22.3 60.9 63.1
CoTracker [cotrackerv1] 55.1 48.9 18.8 18.1 50.8 46.1
CoTracker3 [cotrackerv3] 72.7 71.3 29.6 29.1 65.6 70.6
CoTracker3 [cotrackerv3] 66.3 66.3 25.8 24.9 59.2 63.6
TAPTR [TAPTR] 69.1 59.4 23.7 23.8 61.2 58.0
TAPTR [TAPTR] 63.0 48.9 20.4 19.0 52.3 39.4
TAPIR [doersch2023tapir] 64.2 59.6 21.3 24.6 56.8 50.4
LocoTrack [LocoTrack] 68.7 67.1 25.2 30.2 62.7 69.2
Track-On [Track-On] 69.7 69.7 24.8 25.8 64.5 64.5
[…]
@@
Online Point Tracking. We compare our method to prior works in Table 5 on four benchmarks, including BADJA [badja] and Perception Test [patraucean2023perception] for key point tracking, TAP-Vid [tapvid] for arbitrary point tracking, and our TAG dataset in the ‘query first’ evaluation, which means points appearing in the first frame are used as queries. We report the Percentage of Correct Keypoint-Transfer (PCK-T) for the BADJA benchmark, and Average Jaccard (AJ) for the remaining benchmarks. However, most of the current methods are offline trackers, which process long-temporal window frames or even the entire video to be able to see the future frame, and do not match the online setup required by real applications. For a fair comparison, we modified their input so that there is no future information inside the window, denoted as model, to enable inference online. Experimental results show a substantial decrease in model performance when the input data is switched from offline to online. The comparative analysis reveals the effectiveness of our approach on keypoint tracking benchmarks, which surpasses competing models. Furthermore, although our model is trained on keypoint datasets, it demonstrates generalization capability on arbitrary point tracking datasets.
@@ 6.2 Exploration Studies
Study on data engine. To validate the effectiveness of our data engine, we evaluated the performance when trained on different phases of our TAG dataset, as shown in Table 6. The results demonstrate that our proposed dataset enhances the performance on other datasets, indicating high diversity and generalizability. After training with more data from phases ② and ③, the performance is further improved across all three tasks, demonstrating the effectiveness of the supplementary data provided by our data engine.
@@ 6.3 Visualization
To further illustrate the varying requirements for memory representation of targets at different granularities, we visualize the memory-related outputs for the three tasks under both task-adaptive memory mechanism (left) and full parameter sharing (right), as shown in Fig. 6. Firstly, we observe that even under different training settings, the memory features for the same task remain highly similar, indicating that different granularities have distinct memory requirements. Secondly, under the memory-related decoupled training setting with our task-adaptive memory mechanism, both memory attention and mask output align more closely with the task outputs compared to full parameter sharing, highlighting the necessity of the decoupled design. Finally, we find that under the full parameter sharing setting, the mask output for point tracking does not exhibit a Gaussian pattern, leading to incorrect predictions, while the training method with a decoupled design does not exhibit this error. This demonstrates that the decoupled design effectively preserves the specific needs of different tasks.
@@ 8.2 Training and Inference Details
Training is performed jointly on data of the three tasks. In addition to our Tracking-Any-Granularity dataset, we used DAVIS-17 [davis17], YoutubeVOS-19 [youtube-vos] and MOSE [MOSE] for the mask task, LaSOT [lasot], GOT-10k [got10k], TrackingNet [TrackingNet] and COCO [coco] for the box task, and TAP-Vid Kinetics [tapvid], PointOdyssey [zheng2023pointodyssey], and PerceptionTest [patraucean2023perception] for the point task. To enable the model to be simultaneously capable of all three tasks and to optimize training efficiency, we adopt the strategy of alternating between the three tasks. Specifically, we implement parallelisation by sampling a whole batch at each step of training, which is entirely derived from the data of a particular task. The sampling probability is set to 1:4:5 to balance the performance of the three tasks.
@@
In addition, if interactive input is used as initial prompt, we select up to 2 frames as corrective frames to add corrective clicks on them: after predicting the selected frame, we sample a positive point from the false positive region between the prediction and ground truth or a negative point from the false negative region as a corrective point, and use it as additional prompt to get a new prediction along with all previous cumulative prompt from that frame. This operation is repeated until 7 corrective points have been added. In addition, if the box tracking task uses the box format to compute the regional differences between the prediction and the GT, there is an overwhelming problem that the sampled corrective clicks may fall at the boundaries of the box instead of inside the target, which is contrary to the actual interaction. Therefore, we choose to compute the difference in mask format, and use SAM 2 and sam-hq [sam_hq] with box annotations to obtain the pseudo-GT mask on SOT datasets because of its good segmentation ability.
Losses and optimization. Following the mask tracking task in SAM 2, we adopt the linear combination of focal loss $\mathcal{L}_{mask}^{focal}$ and dice loss $\mathcal{L}_{mask}^{dice}$ for the mask prediction, L1 loss for the IoU prediction $\mathcal{L}_{L1}^{IoU}$, and cross-entropy loss for object occlusion prediction $\mathcal{L}_{CE}^{obj}$. During the box tracking task, we adopt the corner head to predict the bounding boxes and add additional ciou loss [zheng2020diouciou] and L1 loss to supervise the box prediction. As for the point tracking task, we select the highest probability position from the mask prediction as point prediction, and use soft argmax [softargmax] during training for making the process derivable instead of the undifferentiable argmax function. Beyond the loss on mask in the form of Gaussian map, we add an L1 loss between the prediction and ground-truth point to directly optimize the distance and accuracy of the points. For multi-prediction scenario, we only supervise the task predictions (masks, boxes, and points) with the lowest loss, which is a combination of $\mathcal{L}^{mask}$, $\mathcal{L}^{box}$ and $\mathcal{L}^{point}$, but supervise the IoU predictions of all task predictions to learn to synchronise the quality of predictions. Furthermore, if the target is missing in some frames due to disappearance or cropping, we do not supervise the task predictions or iou predictions on them in all three tasks, but always supervise the occlusion prediction from an MLP head, no matter if the ground-truth exists or not. In summary, the supervision losses for the three tasks can be written as:
@@ 9.2 Annotation Pipeline
3) Fine Annotation. To reduce annotator workload and improve efficiency, we use SAM [kirillov2023seganysam1] to generate rough masks based on the coarse annotations (points and boxes). Then, annotators refine these masks with the following requirements:
@@ 10.1 Performance Comparison
Evaluation metrics. In video object segmentation task, we use standard metrics [vos-benchmark] in most benchmarks: Jaccard index $\mathcal{J}$, contour accuracy $\mathcal{F}$, and their average $\mathcal{J\&F}$. In the YouTubeVOS benchmark, $\mathcal{J}$ and $\mathcal{F}$ are computed for ”seen” and ”unseen” categories separately. $\mathcal{G}$ is the averaged $\mathcal{J\&F}$ for both seen and unseen classes. In LVOS benchmark, the first densely annotated long-term VOS dataset with high-quality annotations, it introduces the standard deviation $\mathcal{V}$ of the average score of $\mathcal{J}$ and $\mathcal{F}$ to assess the temporal stability of VOS models. In VOST benchmark, which focuses on segmenting objects as they undergo complex transformations, it additionally reports $\mathcal{J}_{tr}$ for the last 25% of the frames in a sequence to show the robustness after the transformation has been mostly completed. The BURST benchmark is evaluated with Higher Order Tracking Accuracy (HOTA) [hota] as a good balance between measuring frame-level detection and temporal association accuracy. In single object tracking task, we evaluate performance with Area Under the Curve (AUC), normalized precision (${P}_{Norm}$) and precision (P{P}) to measure the average accuracy of center, size and scale between the prediction and labeled groundtruth bounding boxes of all the frames for most benchmarks. For the GOT-10k benchmark, we choose the average overlap (AO) and success rate (SR) as indicators. The former AO denotes the average of overlaps between all groundtruth and estimated bounding boxes, while the SR measures the percentage of successfully tracked frames where the overlaps exceed a threshold (e.g., 0.5). In point tracking task, we report Occlusion Accuracy (OA) and Average Jaccard (AJ) for TAP-Vid [tapvid], Perception Test [patraucean2023perception], and our dataset. As for the BADJA benchmark, we adopt the Percentage of Correct Keypoint-Transfer (PCK-T). We measure these benchmarks in the ‘query first’ evaluation, which means points appearing in the first frame are used as queries.
@@
Unicorn [Unicorn] 69.2 65.2 73.2 83.0 86.4 82.2
UNINEXT-H [UNINEXT] 81.8 77.7 85.8 85.4 89.0 86.4
MITS [MITS] 84.9 82.0 87.7 83.4 88.9 84.6
OmniTracker-L [OmniTracker] 71.0 66.8 75.2 83.4 86.7 82.3
[…]
@@
Comparison with TAG model. We found that there exists a work, TAG [harley2024tag], that shares the same objective as ours, which is to achieve tracking of mask, box, and point with a unified model. To highlight the contribution of our work, we provide a detailed comparison between the two approaches. Firstly, TAG is an offline tracking model that processes multiple frames as a clip simultaneously, which not only differs from the current mainstream online frame-by-frame tracking pipeline but also leads to information leakage from future frames and is only applicable to pre-recorded videos. In contrast, our model ensures that the current frame only receives information from the past frames, making it suitable for video streams. Secondly, in the way of the prompt construction, TAG simply converts point coordinates into a {0, 1} mask in the point task, providing limited target information. Our method combines point coordinates with a (0, 1) Gaussian mask: the former provides precise locations, while the latter highlights the target point in mask form, maintaining consistency with output from MaskDecoder and input for MemoryEncoder, thereby enhancing expressiveness. For the box task, TAG converts the box into a square mask, which causes confusion between the target region and the background, affecting the accuracy of the target information. Third, TAG is trained only on public datasets, which limits the scale of the training dataset. In contrast, we construct a Data Engine that enables both model training and dataset annotation expansion, ultimately resulting in a large-scale dataset with three types of granularity annotations and a well-trained model. Most important of all, as an offline approach, the TAG model primarily focuses on how to jointly encode targets of varying granularity. When processing the next clip, the prompt remains in an original, unmodeled form, lacking rich target representation (e.g., mask, point, or box). Meanwhile, due to the lack of judgment of predictions, the next clip must adopt the prediction of the last frame in the previous clip as a prompt, even if it may be unreliable, which leads to error accumulation and makes it difficult to handle common challenges such as temporary target disappearance. In contrast, our method follows the online setting. The core challenges lie not only in multi-granularity prompts encoding, but also in how to transform predictions into memory representations to guide subsequent frames. Compared to original prompts, memory offers richer target features that improve the stability and accuracy. Leveraging both selective capability and memory diversity, the mechanism compensates for potential errors in individual predictions, effectively improving stability. To make this key component compatible with multi-granularity tracking, we introduced a Task-Adaptive Memory mechanism, one of the major contributions, to unify the predictions of varying granularity into memory representations. In summary, the distinction is not merely at the task level, but it directly impacts motivation and core innovations. Additionally, we present a performance comparison on the three tasks. As shown in the Table 14, our model significantly outperforms the TAG model on all three tasks, demonstrating the superior performance of our model.